# Unveiling the inconsistency of the Proca theory with non-minimal coupling to gravity

### v3

@@ Abstract
We study the degrees of freedom of the Proca theory, non-minimally coupled to gravity. In the Minkowski background, this theory propagates five degrees of freedom – a massive longitudinal mode, two massive vector ones, and two massless tensor modes. At first sight, the non-linear coupling between the metric perturbations and the vector field indicates that both longitudinal and tensor modes become strongly coupled, at the same scale. This would imply that no matter how small the photon mass is if non-minimal coupling is taken into account, gravitational waves would necessarily be strongly coupled. We show that the way out of this inconsistency is through the introduction of the disformal coupling to the metric perturbations that resemble the vector-type disformal transformations. This way, the unphysical coupling between the two types of modes can be avoided, rendering the model consistent. As a result, we show that only the longitudinal modes enter a strong coupling regime, while both tensor and transverse modes remain weakly coupled at all scales up to the Planck length. Finally, we show how this cures recently reported runaway modes in curved spacetime.
@@ 1 Introduction
Intuitively, we can think of the Proca theory as a small modification of Maxwell’s electrodynamics, that amounts to adding a mass term by hand. Then, per the principle of continuity, one might expect that in the limit when this mass is set to zero, all observables of the modified theory should smoothly go back to the original ones [27]. In general, however, it took several decades to show that this principle holds in massive gauge theories. The two particularly interesting examples are massive Yang-Mills theory (mYM) with mass added by hand, and massive linearized gravity [28, 29]. Studying the two with conventional methods could lead to the conclusion that they might not be realized in nature. In the case of mYM theory, when calculating the corrections to the propagator of the non-Abelian vector field it was shown that the perturbative series are singular in mass [30, 31, 32, 33]. For the mGR case, the situation was even more drastic. It yielded predictions that were different from those of Einstein’s relativity and remained such in the massless limit. As pointed out in [30, 34], the occurrence of this perturbative discontinuity in mass was due to the appearance of the extra degrees of freedom (dof), the longitudinal modes, that failed to decouple from the remaining dof in the massless limit. However, as pointed out in [35] in the context of massive gravity, once the non-linear terms become of the same order as the linear ones, the perturbation theory for the longitudinal mode breaks down. Subsequently, on smaller distances than this Vainshtein radius, the longitudinal modes are strongly coupled and one obtains predictions of linearized gravity up to small corrections [36, 37]. The possibility of the smooth massless limit for the mYM was conjectured in [38] as well, and later shown in [39]. The Vainsthein mechanism – mechanism which is responsible for this behavior of the modes absent in massless theories was also generalized to other field theories [40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51]. Additionally, it also holds for the Proca theory, in the presence of the quartic and cubic self-interactions, and generalized Proca theories [52, 53, 50, 54].
@@
The main question that we aim to answer in this paper is – Does Proca theory non-minimally coupled to gravity have strong coupling, and if yes, how strongly coupled the theory is? Based on the results in massive gauge theories, there are three probable outcomes: First, the theory could be already well-behaved in the massless limit, despite the presence of the longitudinal modes. In contrast, the second two possibilities contain perturbative series that are singular in mass, and consequently, which have a strong coupling scale. However, either only the longitudinal mode becomes strongly coupled, or the remaining modes of the theory become strongly coupled as well. From the perspective of self-interacting theory, the first case is natural. In cubic or quartic self-interactions, the transverse modes, dof which are also present in Maxwell’s electrodynamics, remain weakly coupled even after the perturbation theory for the longitudinal modes breaks down. The second outcome, in which the transverse modes, or tensor modes – modes of the metric perturbations – become strongly coupled, however, would be unphysical. It would indicate that in the presence of non-minimal coupling, one could either not be able to recover electromagnetism or discover gravitational waves.
@@
By studying the Proca theory, non-minimally coupled to gravity in a flat background, and taking into account the metric perturbations around it, we will show that the last case takes place – both tensor and longitudinal modes become strongly coupled, yielding a physically inconsistent result. Moreover, we will show that the way to resolve this ambiguity lies in the introduction of the disformal coupling to the metric perturbations simultaneously. Such disformal transformations, which modify the original metric by adding a bilinear term in the external fields or their field-strength have gained a lot of attention in cosmology, and construction of new theories of gravity [55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74]. In addition, we will show that the production of runaway modes – modes with negative frequency in curved background – recently reported in [26], can be cured with the same transformations.
@@
The paper is organized in the following way. First, we will review the free Proca theory and linearized gravity in a flat space-time background, and identify potential sources of the strong coupling. Then, we will take into account the non-linear terms, and analyze each type of the non-minimal coupling separately. This will be followed by the resolution of the problem that appears for the tensor modes, and subsequently the resolution of the runaway in the curved space-time. Finally, we will conclude with a discussion of the results and future prospects of the theory.
@@ 2 The basics of linearized gravity and Proca theory
The goal of this section is to analyze the dof of this theory, starting with the flat background. For this, let us perturb the metric in the following way:
@@ 3 Investigating the Strong Coupling
In the previous section, we have seen that each type of mode behaves differently. In contrast to the transverse modes, which are only dependent on the length scale, the amplitude of the quantum fluctuations for the tensor modes is inversely proportional to the Planck mass. Most importantly, the amplitude of quantum fluctuations for the longitudinal modes is singular in the mass of the vector field. Due to this behavior, in the presence of self-interactions such as $\left(A_{\mu}A^{\mu}\right)^{2}$, the longitudinal modes will necessarily enter a strong coupling regime [50] (see also [39] for the non-Abelian generalization).
In this section, we will investigate if the longitudinal modes of the non-minimally coupled theory, described in (1) become strongly coupled as well. For this, let us go beyond the linear analysis. By expanding the metric perturbations around the flat background as in the previous chapter, we find the following Lagrangian density:
@@
In the following analysis, we will decompose the metric and the vector field according to the spatial rotations according to (5) and 11) respectively, and apply the conformal gauge $E=B=0$, together with $F_{i}=0$.
@@ 3.1 $\beta$ coupling case
$\begin{split}\mathcal{L}_{EH}\sim&M_{pl}^{2}\left(2\phi\Delta\psi-3\dot{\psi}\dot{\psi}-\psi\Delta\psi\right)-\frac{M_{pl}^{2}}{4}S_{i}\Delta S_{i}-\frac{M_{pl}^{2}}{8}\partial_{\mu}h_{ij}^{T}\partial^{\mu}h_{ij}^{T}\\ &-\frac{M_{pl}^{2}}{4}\left(\frac{1}{2}h_{ij}^{T}h_{ik}^{T}h_{jk,\mu}^{T,\mu}+\frac{1}{2}h_{ij}^{T}h_{kl}^{T}h_{kl,ij}^{T}-h_{ij}^{T}h_{il,k}^{T}h_{jk,l}^{T}\right),\\ \\ \mathcal{L}_{P}=&\frac{1}{2}\left[A_{0}\left(-\Delta+m^{2}\right)A_{0}+2A_{0}\Delta\dot{\chi}-\left(\dot{\chi}\Delta\dot{\chi}-m^{2}\chi\Delta\chi\right)\right.\\ &\left.+\left(\dot{A_{i}^{T}}\dot{A_{i}^{T}}-A_{i,j}^{T}A_{i,j}^{T}-m^{2}A_{i}^{T}A_{i}^{T}\right)\right]\\ &+\psi\left[\left(A_{j,i}^{T}-A_{i,j}^{T}\right)\left(A_{j,i}^{T}-A_{i,j}^{T}\right)-\left(\dot{A}_{i}-A_{0,i}\right)\left(\dot{A}_{i}-A_{0,i}\right)+m^{2}A_{i}A_{i}\right]\\ &+\frac{1}{2}(3\psi-\phi)\left[\dot{A}_{i}\dot{A}_{i}-A_{j,i}A_{j,i}+A_{j,i}A_{i,j}+A_{0,i}A_{0,i}-2A_{0,i}\dot{A}_{i}+m^{2}A_{0}^{2}-m^{2}A_{i}A_{i}\right]\\ &+\phi\left[\left(\dot{A}_{i}-A_{0,i}\right)\left(\dot{A}_{i}-A_{0,i}\right)+m^{2}A_{0}^{2}\right]-S_{i}\left[\left(\dot{A}_{j}-A_{0,j}\right)\left(A^{T}_{j,i}-A^{T}_{i,j}\right)+m^{2}A_{0}A_{i}\right]\\ &+\frac{1}{2}h_{ij}^{T}\left[\left(A_{k,i}^{T}-A_{i,k}^{T}\right)\left(A_{k,j}^{T}-A_{j,k}^{T}\right)-\left(A_{0,i}-\dot{A}_{i}\right)\left(A_{0,j}-\dot{A}_{j}\right)+m^{2}A_{i}A_{j}\right]\\ \\ \mathcal{L}_{\beta int}=&\frac{\beta}{2}\left[A_{0}^{2}\left(\Delta\phi+3\ddot{\psi}\right)-A_{0}A_{i}\left(\Delta S_{i}+4\dot{\psi}_{,i}\right)\right]\\ &+\frac{\beta}{4}A_{i}A_{j}\left(2\psi_{,ij}+2\Box\psi\delta_{ij}-2\phi_{,ij}+\dot{S}_{j,i}+\dot{S}_{i,j}+\Box h_{ij}^{T}\right)\end{split}$
@@
The tensor modes, however, indicate strong coupling, that appears sooner than the Planck mass is reached. Their first-order corrections can be estimated as:
$h_{ij}^{(1)}\sim\frac{\beta}{M^{2}_{pl}m^{2}L^{4}}.$ (38)
Therefore, they become of the same order as the linear term $h_{ij}^{T(0)}\sim\frac{1}{M_{pl}L}$ at the following scale:
$L_{\beta str}\sim\left(\frac{\beta}{M_{pl}m^{2}}\right)^{1/3}.$ (39)
For the longitudinal modes, this means that the first term on the left-hand side of (28) is no longer vanishing. Once the scale $L_{\beta str}$ is reached, their first-order corrections should be evaluated as:
@@
Therefore, in this section, we have found that both tensor modes and longitudinal modes become strongly coupled once the scale reaches $L_{\beta str}$. However, this is an unphysical result. It would imply that no matter how small the mass of the photon is, gravitational waves would be strongly coupled if non-minimal coupling with the Ricci tensor is taken into account.
In the following section, we will first study the other non-minimal term, and then show how to avoid this inconsistency.
@@ 3.2 $\alpha$ coupling case
Clearly, the term on the right-hand side is a non-vanishing one even within the perturbation theory. Thus, it yields a strong coupling for the longitudinal mode, with the scale given by:
@@
We can notice that this scale arrives in the same manner as the strong coupling scale of the $\beta$ case, with only the coupling interchanged. However, while for the $\beta$ case we found that both tensor and longitudinal modes become strongly coupled, here only the longitudinal modes enter the strong coupling regime. This is a standard property of Proca theory and is thus not as problematic. Moreover, due to the scaling of the remaining modes – the vector and tensor ones – we can be sure that this will be the most dominant term.
@@ 4 How to cure the tensor modes?
In the previous sections, we have studied the $\alpha$ and $\beta$ models, which involved the coupling of the vector field with the Ricci scalar and the Ricci tensor respectively. In the case of the $\alpha$ coupling, we have found that only the longitudinal modes become strongly coupled, due to the quartic self-interaction that appears after one resolves the constraints. In contrast, the coupling with the Ricci tensor proved to be troubling. There, we have found that both tensor and longitudinal modes become strongly coupled, which is a physically unacceptable result.
Therefore, one might accept that this is the case, and exclude such interaction in the further models. In this work, we will, however, try to resolve this puzzle in another way. Notably, a similar problem can be encountered in massive Yang-Mills theory, with mass added by hand [39]. If one decomposes the spatial part of this vector field only linearly, in a way done in this work, one would necessarily encounter the strong coupling of the transverse modes. However, by correcting the decomposition non-linearly, one can obtain that only the longitudinal modes become strongly coupled.
Let us apply the procedure found in [39] to cure the strong coupling of the transverse modes to this case as well. We can notice that at the strong coupling scale $L_{\beta str}$, both terms are of the same order in the equation of motion for the tensor modes:
@@
However, the form of this equation suggests that the tensor modes should be transformed:
@@
This form of the introduction of new couplings suggests that writing the metric perturbations in a manifestly covariant way, also known as the disformal transformation, as
$g_{\mu\nu}=\eta_{\mu\nu}+h_{\mu\nu}-\frac{\beta}{M_{pl}^{2}}A_{\mu}A_{\nu}$ (55)
could remove the problematic coupling between the tensor and longitudinal modes. In this section, we will explore if this will indeed be the case. In particular, the goal of the following analysis is to show that with the above-modified decomposition, both tensor and transverse modes will be weakly coupled at all scales up to the Planck length, while the longitudinal modes will become strongly coupled once the scale $L_{\beta str}$ is reached.
In order to show this, let us study the longitudinal modes in quartic order. With (55), we find the following Lagrangian density:

### v4

@@ Abstract
We study the degrees of freedom of the Proca theory, non-minimally coupled to gravity. In the Minkowski background, this theory propagates five degrees of freedom – a massive longitudinal mode, two massive vector ones, and two massless tensor modes. At first sight, the non-linear coupling between the metric perturbations and the vector field indicates that both longitudinal and tensor modes become strongly coupled, at the same scale. This would imply that no matter how small the photon mass is if non-minimal coupling is taken into account, gravitational waves would necessarily be strongly coupled. We show that the way out of this inconsistency is through the introduction of the disformal coupling to the metric perturbations that resemble the vector-type disformal transformations. This way, the unphysical coupling between the two types of modes can be avoided, rendering the model consistent. As a result, we show that only the longitudinal modes enter a strong coupling regime, while both tensor and transverse modes remain weakly coupled at all scales up to the Planck length. Finally, using the same form of the disformal transformation, we introduce a disformal frame in which the recently reported runaway modes are absent.
@@ 1 Introduction
A natural question that arises when studying massive gauge theories is – How do they behave when the mass is set to zero? According to the principle of continuity, in such a limit, all observables should match those of the corresponding massless theory [27]. Interestingly, however, the first insights into such question posed non-trivial challenges, originally noticed in massive Yang-Mills theory (mYM) with mass added by hand, and massive linearized gravity [28, 29]. Studying the two with conventional methods could lead to the conclusion that they might not be realized in nature. In the case of mYM theory, when calculating the corrections to the propagator of the non-Abelian vector field it was shown that the perturbative series are singular in mass [30, 31, 32, 33]. For the mGR case, the situation was even more drastic. It yielded predictions that were different from those of Einstein’s relativity and remained such in the massless limit.
@@
As pointed out in [30, 34], the occurrence of this perturbative discontinuity in mass was due to the appearance of the extra degrees of freedom (dof), the longitudinal modes, that failed to decouple from the remaining dof in the massless limit. However, in [35], it was noticed in the context of massive gravity that once the non-linear terms become of the same order as the linear ones, the perturbation theory for the longitudinal mode breaks down. Subsequently, on smaller distances than this Vainshtein radius, the longitudinal modes are strongly coupled, and one obtains predictions of linearized gravity up to small corrections [36, 37]. The possibility of the smooth massless limit for the mYM was conjectured in [38] as well, and later shown in [39]. The Vainsthein mechanism – mechanism which is responsible for this behavior of the modes absent in massless theories was also generalized to other field theories [40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54].
Similarly to the mGR and mYM theories, Proca theory with cubic and quartic self-interactions obeys the Vainshtein mechanism as well - the longitudinal mode that is absent in the massless theory indicates a discontinuity in the perturbative series, becomes strongly coupled once this series breaks down and decouples from the remaining dof beyond such scales, which in turn remain weakly coupled [41, 50]. In this paper, we will also explore the massless limit of the Proca theory, now focusing on its non-minimal coupling to gravity, through the Ricci scalar and the Ricci tensor. In particular, motivated by the previous studies on Proca theory, and massive gauge theories in general, we will aim to answer – Does Proca theory with non-minimal coupling to gravity have strong coupling in flat spacetime, and if yes, how strongly coupled the theory is?
@@
Based on the results of massive gauge theories, there are three possible outcomes: First, the theory could be already well-behaved in the massless limit, despite the presence of the longitudinal modes. The theory might likewise contain perturbative series that are singular in mass, which would consequently imply strong coupling. As the longitudinal modes are absent in the massless theory, it would be natural to expect that they become strongly coupled, with the rest of the modes remaining in the weakly coupled regime. This has been shown in the presence of cubic or quartic self-interactions. Another possibility is that the remaining modes enter the strong coupling regime as well. This outcome, in which the transverse modes, or tensor modes – modes of the metric perturbations – become strongly coupled, however, would be unphysical. It would indicate that in the presence of non-minimal coupling, one could either not be able to recover electromagnetism or discover gravitational waves.
By studying the Proca theory, non-minimally coupled to gravity in a flat background, and taking into account the metric perturbations around it, we will show that, surprisingly, the last case takes place. Similarly to mYM theory, both types of coupling will indicate a perturbative series that is singular in mass. At the same time, following the close analogy to the Vainshtein mechanism, due to the non-linear terms, the perturbation theory will break down at the corresponding scales. As we will see, while the Ricci scalar coupling introduces a natural strong coupling that involves only the longitudinal modes, the coupling of the Proca field to the Ricci tensor introduces a strong coupling in both longitudinal and tensor modes, thus yielding in a physically inconsistent result.
@@
Moreover, we will show that the way to resolve this ambiguity lies in the introduction of the disformal coupling to the metric perturbations simultaneously. Such disformal transformations, which modify the original metric by adding a bilinear term in the external fields or their field strength have gained a lot of attention in cosmology, and construction of new theories of gravity [55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74]. As we will see, in flat spacetime, the background value for the Ricci tensor is vanishing, so the form of the disformal coupling necessary to remove the tensor mode ambiguity will be possible to deduct by studying the form of the coupling already on the level of action. In this paper, we will also infer and confirm its form by closely following the analogy with the mYM theory.
Recently, in [26] it was reported that the Proca theory with non-minimal coupling to gravity suffers from another problem in the curved background – the production of runaway modes which have a negative frequency. In this work, we will show that the origin of this instability lies due to the coupling of the vector field with the Ricci tensor. While generically the instability might persist, we will show that in the disformal frame, defined by the same form of the disformal transformation as in the flat background, and in which the matter is minimally coupled to gravity, the runaway modes are absent.
@@
The paper is organized in the following way: In section 2, we will review the free Proca theory and linearized gravity in a flat space-time background. In section 3, we will study the massless limit of the Proca theory with non-minimal couplings to gravity. By treating the Ricci scalar and Ricci tensor couplings separately, we will show the appearance of unphysical behavior of the tensor and longitudinal modes in the case of the latter interaction. Following the close analogy with the mYM theory, we will introduce the resolution to this puzzle in section 4. This will be followed by the resolution of the runaway in the curved space-time in section 5. Finally, we will conclude with a discussion of the results and prospects of the theory.
@@ 2 The basics of linearized gravity and Proca theory
The goal of this section is to analyze the degrees of freedom dof of this theory, starting with the flat background. For this, let us perturb the metric in the following way:
@@ 3 Investigating the Strong Coupling
In the previous section, we have seen that each type of mode behaves differently. In contrast to the transverse modes, which are only dependent on the length scale, the amplitude of the quantum fluctuations for the tensor modes is inversely proportional to the Planck mass. Most importantly, the amplitude of quantum fluctuations for the longitudinal modes is singular in the mass of the vector field.
The main focus of this section will be to investigate the massless limit of the Proca theory, with non-minimal coupling to the Ricci tensor and Ricci scalar given in (1). Due to their amplitude of quantum fluctuations, we might expect that the longitudinal modes might give rise to the perturbative series that is singular in mass. This is the case in the presence of self-interactions such as $\left(A_{\mu}A^{\mu}\right)^{2}$, where the longitudinal modes consequently enter a strong coupling regime [50] once the non-linear terms become of the same order as the linear ones (see also [39] for the non-Abelian generalization).
To clearly see the structure of each interaction, we will consider the Ricci scalar and the Ricci tensor type couplings separately, without any loss of generality. To find the behavior of the leading order corrections, we will follow the procedure introduced [49], going beyond the linear analysis and estimating the corrections to each of the modes based on the minimal amplitude of the quantum fluctuations.
By expanding the metric perturbations around the flat background as in the previous chapter, we find the following Lagrangian density:
@@
In the following analysis, we will decompose the metric and the vector field according to the spatial rotations given in (5) and 54) respectively, and apply the conformal gauge $E=B=0$, together with $F_{i}=0$.
@@ 3.1 $\beta$ coupling case
$\begin{split}\mathcal{L}_{EH}\sim&M_{pl}^{2}\left(2\phi\Delta\psi-3\dot{\psi}\dot{\psi}-\psi\Delta\psi\right)-\frac{M_{pl}^{2}}{4}S_{i}\Delta S_{i}-\frac{M_{pl}^{2}}{8}\partial_{\mu}h_{ij}^{T}\partial^{\mu}h_{ij}^{T}\\ &-\frac{M_{pl}^{2}}{4}\left(\frac{1}{2}h_{ij}^{T}h_{ik}^{T}h_{jk,\mu}^{T,\mu}+\frac{1}{2}h_{ij}^{T}h_{kl}^{T}h_{kl,ij}^{T}-h_{ij}^{T}h_{il,k}^{T}h_{jk,l}^{T}\right),\end{split}$
$\begin{split}\mathcal{L}_{P}=&\frac{1}{2}\left[A_{0}\left(-\Delta+m^{2}\right)A_{0}+2A_{0}\Delta\dot{\chi}-\left(\dot{\chi}\Delta\dot{\chi}-m^{2}\chi\Delta\chi\right)\right.\\ &\left.+\left(\dot{A_{i}^{T}}\dot{A_{i}^{T}}-A_{i,j}^{T}A_{i,j}^{T}-m^{2}A_{i}^{T}A_{i}^{T}\right)\right]\\ &+\psi\left[\left(A_{j,i}^{T}-A_{i,j}^{T}\right)\left(A_{j,i}^{T}-A_{i,j}^{T}\right)-\left(\dot{A}_{i}-A_{0,i}\right)\left(\dot{A}_{i}-A_{0,i}\right)+m^{2}A_{i}A_{i}\right]\\ &+\frac{1}{2}(3\psi-\phi)\left[\dot{A}_{i}\dot{A}_{i}-A_{j,i}A_{j,i}+A_{j,i}A_{i,j}+A_{0,i}A_{0,i}-2A_{0,i}\dot{A}_{i}+m^{2}A_{0}^{2}-m^{2}A_{i}A_{i}\right]\\ &+\phi\left[\left(\dot{A}_{i}-A_{0,i}\right)\left(\dot{A}_{i}-A_{0,i}\right)+m^{2}A_{0}^{2}\right]-S_{i}\left[\left(\dot{A}_{j}-A_{0,j}\right)\left(A^{T}_{j,i}-A^{T}_{i,j}\right)+m^{2}A_{0}A_{i}\right]\\ &+\frac{1}{2}h_{ij}^{T}\left[\left(A_{k,i}^{T}-A_{i,k}^{T}\right)\left(A_{k,j}^{T}-A_{j,k}^{T}\right)-\left(A_{0,i}-\dot{A}_{i}\right)\left(A_{0,j}-\dot{A}_{j}\right)+m^{2}A_{i}A_{j}\right]\\ \\ \mathcal{L}_{\beta int}=&\frac{\beta}{2}\left[A_{0}^{2}\left(\Delta\phi+3\ddot{\psi}\right)-A_{0}A_{i}\left(\Delta S_{i}+4\dot{\psi}_{,i}\right)\right]\\ &+\frac{\beta}{4}A_{i}A_{j}\left(2\psi_{,ij}+2\Box\psi\delta_{ij}-2\phi_{,ij}+\dot{S}_{j,i}+\dot{S}_{i,j}+\Box h_{ij}^{T}\right)\end{split}$
@@
In contrast to the transverse and the longitudinal modes, we can notice that the first-order corrections, which can be estimated as:
$h_{ij}^{(1)}\sim\frac{\beta}{M^{2}_{pl}m^{2}L^{4}},$ (38)
are singular in mass. Once they become of the same order as the linear term $h_{ij}^{T(0)}\sim\frac{1}{M_{pl}L}$, the tensor modes enter a strong coupling regime. This corresponds to the following length scale:
$L_{\beta str}\sim\left(\frac{\beta}{M_{pl}m^{2}}\right)^{1/3},$ (39)
appearing sooner than the Planck length is reached.
As the perturbation theory is no longer applicable at scales corresponding to the previous strong coupling scale, the first term on the left-hand side of the equation for the longitudinal modes (28) is no longer vanishing. Once the scale $L_{\beta str}$ is reached, their first-order corrections should be evaluated rather as:
@@
When studying Proca theory, with interactions that are perturbatively singular in the limit when the mass goes to zero, it is natural to expect that the longitudinal mode becomes strongly coupled, as this is the source of the singular behavior, and is absent in the massless theory. Yet, in this subsection, we find that due to the coupling with the Ricci tensor, not only the longitudinal mode becomes strongly coupled, but the tensor modes as well, once the scale $L_{\beta str}$ is reached. However, this is an unphysical result. It would imply that no matter how small the mass of the photon is, gravitational waves would be strongly coupled if non-minimal coupling with the Ricci tensor is taken into account.
Let us now study the coupling with the Ricci scalar, and evaluate if the same effect takes place.
@@ 3.2 $\alpha$ coupling case
The term on the right-hand side is not vanishing even within the perturbation theory. By taking into account the minimal amplitude of quantum fluctuations for the longitudinal mode, we can notice that it is also singular in mass. It will become of the same order as the linear term $\chi^{(0)}$ at the strong coupling scale:
@@
We can notice that this scale arrives in the same manner as the strong coupling scale of the $\beta$ case, with only the coupling interchanged. However, while for the $\beta$ case we found that both tensor and longitudinal modes become strongly coupled, here only the longitudinal modes enter the strong coupling regime. This is also to be expected – the longitudinal modes are absent in the massless theory, while the tensor and transverse modes remain there. Moreover, due to the scaling of the remaining modes – the vector and tensor ones – we can be sure that this will be the most dominant term.
@@ 4 How to cure the tensor modes?
In the previous sections, we have studied the $\alpha$ and $\beta$ models, which involved the coupling of the vector field with the Ricci scalar and the Ricci tensor respectively. By writing the theory only in terms of the propagating modes, and by studying their corresponding equations of motion, we have shown that within the perturbation theory, both types of coupling are seemingly singular in mass, and consequently introduce strong coupling scales. In the case of the Ricci scalar, we find that only the longitudinal mode becomes strongly coupled, due to the quartic self-interaction that appears after one resolves all of the constraints. In contrast, the coupling of the vector field with the Ricci tensor indicates that both the longitudinal mode and the tensor mode become strongly coupled, leading us to a physical ambiguity.
On the one hand, one might simply discard the couplings with the Ricci tensor and focus only on the coupling to the Ricci scalar as a viable coupling. In this work, we will, however, try to resolve this puzzle in another way.
Interestingly, a similar problem appears in the case of massive Yang-Mills theory, with mass added by hand [39]. In flat space-time, this theory describes the massive SU(2) vector field $A_{\mu}$ and is given by the action:
$S=\int d^{4}x\left[-\frac{1}{2}\text{Tr}(F_{\mu\nu}F^{\mu\nu})+m^{2}\text{Tr}(A_{\mu}A^{\mu})\right],$ (53)
where
$F_{\mu\nu}=D_{\mu}A_{\nu}-D_{\nu}A_{\mu},\qquad\text{and}\qquad D_{\mu}=\partial_{\mu}+igA_{\mu}$ (54)
are the field strength tensor and the covariant derivative respectively, while $g$ is the coupling constant. Similarly to the Proca theory, the temporal component of the vector field is not propagating. If one finds its constraint, solves it, and substitutes it back to the action, one will find the action only in terms of the transverse and the longitudinal modes. Similarly to (), we can decompose the spatial part of the vector field as:
$A_{i}^{a}=A_{i}^{Ta}+\chi_{,i},\qquad\text{where}\qquad A_{i,i}^{Ta}=.$ (55)
Here, we have expressed the non-Abelian vector field in terms of the generators of the SU(2) group
$A_{\mu}=A_{\mu}^{a}T^{a},\qquad a=1,2,3,$ (56)
where $T^{a}=\frac{\sigma^{a}}{2}$ are the Pauli matrices. However, in the case of the decomposition (55), one will find that the transverse modes enter the strong coupling regime in addition to the strong coupling of the longitudinal modes, in a way that is similar to the tensor modes in the $\beta$ coupling case:
$(\Box+m^{2})A_{i}^{Ta}\sim g\varepsilon^{abc}P_{ij}^{T}\left[\frac{1}{2}\Box\left(\chi^{b}\chi^{c}_{,j}\right)\right],\qquad P_{ij}^{T}=\delta_{ij}-\frac{\partial_{i}\partial_{j}}{\Delta}$ (57)
The right-hand side of the above equation, however, suggests that the introduction of the coupling in the form of such a term could remove the strong coupling of the transverse modes. Notably, by introducing the coupling in the following way,
$\begin{split}A_{i}^{a}=A_{i}^{Ta}+\chi_{,i}^{a}-g\varepsilon^{abc}(\frac{1}{2}\chi_{,i}^{b}\chi^{c}),\end{split}$ (58)
one not only removes the above strong coupling of the transverse modes but also shifts the strong coupling scale. In fact, by studying the new theory with the coupling given by (58), one will still find unphysical results for the transverse modes, now at a higher scale, and should repeat the procedure by introducing an infinite number of couplings, defined in the following way:
$A_{i}=\zeta A_{i}^{T}\zeta^{\dagger}+\frac{i}{g}\zeta_{,i}\zeta^{\dagger}\qquad\text{with}\qquad A_{i,i}^{T}=,$ (59)
where $\zeta$ is the unitary matrix given by
$\zeta=e^{-ig\chi}.$ (60)
In this case, only the longitudinal modes will become strongly coupled, with the strong coupling scale that matches the scale at which the unitarity is violated.
Following the case for the mYM theory, we could introduce additional coupling to the theory such that the strong coupling of the tensor mode disappears. Following the procedure found in [39], we can notice that at the strong coupling scale $L_{\beta str}$, both terms are of the same order in the equation of motion for the tensor modes:
@@
The form of this equation suggests that the tensor modes should be transformed:
@@
This form of the introduction of new couplings, known as the disformal couplings suggests that writing the metric perturbations in a manifestly covariant way as:
$g_{\mu\nu}=\eta_{\mu\nu}+h_{\mu\nu}-\frac{\beta}{M_{pl}^{2}}A_{\mu}A_{\nu},$ (63)
could remove the problematic coupling between the tensor and longitudinal modes. Notably, this form could be also deducted directly from the action. For this, one should note that the Ricci tensor vanishes for the flat background, thus having the dependence only on the perturbations. Therefore, one can trace the appearance of the tensor modes among them, and then by looking at the form of the coupling infer the resulting disformal transformation.
In the remaining part of this section, we will study the disformal theory, and investigate if the disformal coupling removes the strong coupling of the tensor modes. In addition, we will have to make sure if the coupling previously introduced is the only necessary coupling, or similarly to the mYM case one has to introduce multiple ones. Finally, we will verify if the longitudinal modes nevertheless enter the strong coupling regime, and if the resulting strong coupling scale differs from the one that we have initially found. To answer these questions, let us study the longitudinal modes at the quartic order.
With (63), we find the following Lagrangian density:
@@
Therefore, we have found that the introduction of the disformal coupling of the form (63) removes the strong coupling of the tensor modes. At the same time, the strong coupling of the longitudinal modes remains in the theory. Curiously, we should notice that, unlike the mYM case, the strong coupling scale does not change upon the introduction of the disformal coupling. In the following section, we will see the benefits of the introduction of the coupling (63) also for the cosmological background.